# Haptic bilateral teleoperation system for free-hand dental procedures

### v2

@@ IV Experimental validation
To evaluate the performance of the proposed FH-HBTS, we conduct interaction experiments using a phantom. Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary. During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD. We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar, as shown in Figure 6. The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential. For each target, we perform three interaction experiments with active force limitation and three without. With the aim of confirming the impact of the proposed strategy on the safety of interaction, we estimate, from the data recorded during the experiments, the conditional expectation of the interaction forces given a virtual penetration commanded by the operator. The conducted analysis is reported in Section IV-B.
Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right second premolar.
In order to quantitatively assess the operation of the eye-hand coordination, we conduct a free-space experiment which concerns commanding the same spatial motion in the HD space, with different viewing angles. The results of the task execution are reported in Section IV-C.
Finally, we provide an experimental demonstration of the end-to-end system by conducting interaction experiments with the three target teeth in both scenarios, this time emulating the dental scaling procedure. The outcomes of a dental scaling procedure execution, representative of the overall results are shown in Section IV-D.
@@ IV-A Experimental setup
The experimental setup involves the following equipment and materials:
  • a leader control station featuring a 6-degrees-of-freedom (DOF) force-feedback HD [50] and a workstation. The latter is a general purpose computer executing the leader controller loop, and communicating with the FR over an Ethernet network. In such a network, delays and packet losses are negligible, as is the case of an healthcare facility;
  • a FR consisting of the 6-DOF UR10 robotic arm by Universal Robots [51]. It is equipped with an integrated mechanical end-effector for free-hand dental procedures. Its design is discussed in Section IV-A1;
  • a dental phantom, i.e. 3D-printed life-sized model of a human mouth.
IV-A1 Integrated mechanical end-effector
In order to successfully perform any free-hand dental procedure, the operator relies on three basic factors: (a) the procedure-specific clinical tool; (b) the surgical field view, which may require external illumination and tools to show obstructed or poorly lit areas e.g., a mirror; (c) the sense of touch to ensure the right pressure to be exerted on the area of interest, in order to achieve a safe and effective procedure. In this work, we design an integrated mechanical end-effector, shown in Figure 7, that can provide all the tools and sensory capabilities required to perform free-hand dental procedures.
The designed end-effector consists of five main components:
  1. a)
    a 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor;
  2. b)
    a force/torque sensor measuring contact forces. The sensor used is the FT150 6-axis force/torque sensor by Robotiq [52];
  3. c)
    a 3D-printed tool socket supporting the clinical tool and the endoscopic camera. The interface consists of a rigid structure to enable proper transmission of forces from the tool to the sensor, a circular hole, adjustable via a bolt-and-nut mechanism, to accommodate tools of different sizes, and a movable camera-support ring, allowing for adjustments of the view and light beam angle;
  4. d)
    an HD endoscopic camera, equipped with LED illumination to ensure good visibility even in poorly lit areas;
  5. e)
    a clinical tool for performing the procedure of interest. The tool is placed in the supporting socket after being wrapped with friction tape to ensure good mechanical compliance, emulating the skin friction between the operator’s hand and the tool. The tool used for the chosen use case is a scaler.
Figure 7: Integrated mechanical end-effector for free-hand procedures, consisting of: 1) 3D-printed mechanical interface connecting the robot’s flange with the force/torque sensor; 2) force/torque sensor; 3) 3D-printed tool socket supporting the clinical tool and the endoscopic camera; 4) HD endoscopic camera; 5) clinical tool. Visible in the left side of the figure is also the dental phantom used.
IV-A2 Eye-hand coordination controller parametrization
[…]
With regard to the scaling factors, we set $\beta=1$ and ${\boldsymbol{\Lambda}}=diag(0.8,0.8,0.8,1.0,1.0,1.0)$ to achieve high control accuracy on the translational axes without significantly reducing the workspace size.
IV-A3 Admittance controller parametrization
Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set ${\bf K}_{P}=\text{diag}(50000,50000,6000)$, ${\bf K}_{D}=\text{diag}(2500,2500,12000)$ and ${\bf M}_{d}=\text{diag}(50,50,50)$.
Regarding the force limitation strategy, we identify a suitable trade-off between responsiveness and stability by setting $K_{s}=0.06$, and by filtering the exerted forces with a moving average filter of window size $n=500$.
IV-A4 Virtual feedback parametrization
With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations. Therefore, we set ${\bf K}_{h}=\text{diag}(0.1,0.1,0.1)$, ${\bf D}_{h}=\text{diag}(0.5,0.5,0.5)$.
[…]
@@ IV-B Force limitation assessment
[…]
@@
Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B\right)$, and the conditional variance $\text{Var}(A|B)$ following [55]. In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$, in both scenarios. Visual inspection showcases a significant reduction in the expected normal interaction force in Scenario A, given the same virtual penetration commanded.
@@
Figure 8: Comparison between the conditional expectation of the interaction force in Scenario A and Scenario B.
To evaluate the effectiveness of the force limitation strategy, we compare $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ in both scenarios. To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test. Levene’s test shows that the conditional variances are significantly different (p-value $p\approx 0$), indicating that force limitation strategy affects the dispersion of normal force $f_{e}^{\mathcal{R}_{e},\perp}(k)$ with respect to virtual penetration $\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\right\rVert$. Bartlett’s test, which is more sensitive to normal distributions, confirms this observation ($p\approx 0$). This suggests that the difference in variance is not only due to the sampling effect but is a structural feature of the collected data. We compare the conditional expectation using a t-test for independent samples. Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances. The t-test returned a t-statistic value of $t=-6.18$, confirming that the normal force is lower in Scenario A, and $p\approx 0$. Since such value is below the significance threshold $\alpha=0.05$, we can reject the null hypothesis and conclude that the conditional expectations of normal force in the two scenarios are significantly different.
@@ IV-C Eye-hand coordination assessment
In order to perform a formal assessment of eye-hand coordination, we execute a teleoperated task in the free space. The human operator performs a translatory motion with the HD, simultaneously rotating the viewing angle by rotating the HD’s stylus by an angle $\varphi$ about the $\bf z$ axis of $\mathcal{H}_{e}$. The validation is considered completed if the motion transferred to the robot varies as the viewing angle does. Figure 9 shows the $\varphi$ angle, the position of HD’s stylus and the position of the FR, respectively.
(a) $\varphi$ angle about the $z$ axis of $\mathcal{H}_{e}$.
(b) Position of HD’s stylus.
(c) Position of the FR.
@@
[…]
@@
These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field. However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator. Therefore, in the accompanying video [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors. This includes a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration. This material is 51.4 MB in size. we provide a comprehensive qualitative assessment of the latter.
IV-D End-to-end system experimental demonstration
We provide a demonstration of the end-to-end system’s operation in the accompanying video, where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A. In this section, we report the outcomes of the execution of a dental scaling procedure in Scenario A, highlighting two instants of the experiment (Figure 10), and compare them against those obtained in Scenario B (Figure 11). To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the upper right second premolar.
In Figure 10d and Figure 11b, the normal forces exerted by the FR in both scenarios are shown. A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B. The effectiveness of this strategy is made even more evident by comparing Figure 10c and Figure 11a. Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B. Nevertheless, as shown in Figure 10c, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
[…]
(c) Norm of the admittance controller reference
(d) Contact forces
(e) Forces rendered on the HD
Figure 10: Snapshots and outcomes of the dental scaling experiment in Scenario A.
(a) Norm of the admittance controller reference
(b) Contact forces
(c) Forces rendered on the HD
Figure 11: Outcomes of the dental scaling experiment in Scenario B.
Finally, in Figure 10e and Figure 11c the forces rendered on the HD in both scenarios are shown. As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics. Such aspects enhance naturalness for the human operator.
V Discussions and conclusions

### v3

@@ IV Technical validation
To evaluate the performance of the proposed FH-HBTS, we conduct a set of experiments including static force application (Section IV-B), dynamic dental scaling motions (provided in the supplementary materials), and a free-space experiment (Section IV-C). During the experiments, the robot is out of the operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD. We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar. The chosen targets are significant as each requires a different tool orientation, making the use of eye-hand coordination essential.
@@ IV-A Experimental setup
The experimental setup involves the following equipment and materials: (I) a leader control station featuring a 6-degrees-of-freedom (DOF) force-feedback HD [^2] https://www.3dsystems.com/haptics-devices/touch and a workstation: the latter is a general purpose computer executing the leader controller loop, and communicating with the FR over an Ethernet network, where delays and packet losses are negligible; (II) a FR consisting of the 6-DOF UR10 arm [^3] https://www.universal-robots.com/, equipped with the integrated mechanical end-effector of Section III-D, where the force/torque sensor is the Robotiq FT150444https://robotiq.com; (III) a high-strength resin dental model and a 3D-printed dental phantom, both life-sized models of a human mouth. The former is used for the experiments described in this section, while the latter only appears in the accompanying video.
@@ IV-B Force limitation assessment
To assess the effectiveness of the proposed force limitation strategy, we perform $9$ interaction experiments, with an average duration of $\sim$20\text{\,}\mathrm{s}$\$ each, 3 for each target tooth, without force limitation (Scenario A), $9$ more with force limitation (Scenario B) and $9$ with a passivity-based architecture (Scenario C), specifically TDPA-HD, where admittance control replaces Cartesian impedance control adopted in [42]. In the latter scenario, the force limitation is intentionally kept disabled in order to isolate the effects of TDPA-HD. For each scenario, we record the data sequence $\bigl\{a(k),b(k)\bigr\}_{k=1}^{N}$, sampling at a frequency of $125\text{\,}\mathrm{H}\mathrm{z}$, where we set the force norm $\left\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\right\rVert$ as $a(k)\in\mathcal{A}\subseteq\mathbb{R}_{+}$, and the commanded virtual penetration $\bigl\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\bigr\rVert$ as $b(k)\in\mathcal{B}=\{b\in\mathbb{R}|b_{\text{min}}=$0\text{\,}\mathrm{m}$\leq b\leq b_{\text{max}}=$0.007\text{\,}\mathrm{m}$\$\}. We discretize $\mathcal{B}$ with the step $b_{\text{step}}=$0.0001\text{\,}\mathrm{m}$\$, selected based on the repeatability of the FR, while $b_{\text{max}}$ is the maximum virtual penetration observed across all experiments.
@@
By binning data in $\mathcal{B}$, following [56], we estimate the conditional expectation of the exerted force $E\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$, in all three scenarios. Visual inspection of Figure 7 showcases a significant reduction in the expected interaction force in Scenario B, given the same commanded virtual penetration. Notably, TDPA-HD [42] (Scenario C) cannot reduce the interaction force in quasi-static contacts due to the energy being approximately zero.
@@
Figure 7: Comparison between the conditional expectation of the interaction force in all three scenarios.
To evaluate the effectiveness of the force limitation strategy, we compare $E\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ and $\text{Var}\bigl(\lVert{\bf f}_{e}^{\mathcal{R}_{e}}\rVert\big|\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\rVert\bigr)$ in all three scenarios. We assess differences in conditional variances through Levene’s and Bartlett’s tests in the comparisons involving Scenario B, i.e., B–A and B–C. Both tests indicate significantly different variances ($p\approx 0$), showing that the force limitation strategy affects the dispersion of the force norm $\bigl\lVert{\bf f}_{e}^{\mathcal{R}_{e}}(k)\bigr\rVert$ with respect to virtual penetration $\bigl\lVert{\bf p}_{\mathcal{R}{ed}}^{\mathcal{R}_{e}}(k)\bigr\rVert$. Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances. In the B–A comparison, the test yields a t-statistic t=4.52t=-4.52 with $p\approx 0$, confirming a statistically significant force reduction in Scenario B. Beyond statistical significance, the magnitude of this effect is notable: the comparison yields a Cohen’s d=0.63d=-0.63, indicating a significant force reduction when the force limitation strategy is enabled. The corresponding confidence interval [0.91,0.36][-0.91,-0.36] does not include zero, further supporting the robustness and consistency of the observed effect across the explored range of virtual penetration.
An analogous analysis is conducted for the B-C comparison. Also in this case, Welch’s t-test confirms a statistically significant reduction of the normal force with respect to Scenario C ($p\approx 0$), indicating that the force limitation strategy yields lower contact forces than the passivity-based control architecture. The observed effect is comparable in magnitude to that of the B-A comparison (Cohen’s d=0.68d=-0.68 and confidence interval [0.97,0.41][-0.97,-0.41]), further supporting the robustness of the proposed approach across different control architectures.
@@ IV-C Eye-hand coordination assessment
To perform a formal assessment of eye-hand coordination, we execute a teleoperated task in the free space. The operator performs a translatory motion with the HD, simultaneously rotating the viewing angle by rotating the HD’s stylus by an angle $\varphi$ about the $\bf z$ axis of $\mathcal{H}_{e}$. The validation is considered completed if the motion transferred to the robot varies as the viewing angle does. Figure 8 shows the $\varphi$ angle, the position of HD’s stylus and the position of the FR, respectively.
@@
Results showcase that, as long as the angle $\varphi=$0\text{\,}\mathrm{rad}$\$ (from $0\text{\,}\mathrm{s}$ to $10\text{\,}\mathrm{s}$), a motion along the $x$ axis of $\mathcal{H}_{b}$ is mapped into a motion along the $x$ axis of $\mathcal{R}_{b}$. When the angle turns into $\varphi=-\pi/$2\text{\,}\mathrm{rad}$\$, a motion along the $x$ axis of $\mathcal{H}_{b}$ is mapped into a motion along the $z$ axis of $\mathcal{R}_{b}$ (from $20\text{\,}\mathrm{s}$ to $30\text{\,}\mathrm{s}$).
@@
These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field. However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator. Therefore, in the accompanying video we provide a comprehensive qualitative assessment of the latter.
V In-vitro pre-clinical validation
To validate the feasibility of the FH-HBTS in performing free-hand dental procedures, we conduct an in-vitro experimental trial replicating a SRP procedure under more representative clinical conditions. The procedure is carried out on the dental model, coated with an acrylic polymer–based suspension containing titanium dioxide to simulate artificial calculus, using a fully functional ultrasonic scaler activated via a foot pedal at the leader control station.
A total of 20 participants with varying levels of experience in the dental field are recruited: (Group 1) dental students who had performed at least 20 in-vivo dental scaling procedures $(31.6\%)$; (Group 2) dentists with less than 5 years of professional experience $(36.8\%)$; (Group 3) dentists with more than 5 years of professional experience $(31.6\%)$. Before the experimental phase, all participants complete a structured training session aimed at ensuring competent use of the system, during which they practice approaching the dental surface, maintaining stable contact, and performing preliminary calculus removal.
After training, participants perform artificial calculus removal on selected tooth surfaces: lower left central incisor, lower left second molar, lower right first molar, relying on both visual and haptic feedback. For each trial, each surface is uniformly coated with a thin layer of artificial calculus. The procedure is considered successful if the participant is satisfied with the outcome and none of the safety mechanisms described in Section III-B3 are triggered. A single representative trial, illustrating the overall experimental procedure, is provided in the accompanying video.
V-A Data collection and analysis
To comprehensively evaluate system performance, operator experience, and operational safety, a set of objective and subjective outcome measures is collected and analyzed. These include indicators of clinical efficacy, task efficiency, interaction safety, and perceived workload and usability. Differences across participant experience levels are assessed for all outcome measures using Kruskal–Wallis tests.
System efficacy is evaluated by an independent expert dentist, who assesses calculus removal using a six-point ordinal scale ranging from no removal to complete removal. Task efficiency is quantified through the median execution time per dental surface, while interaction safety is assessed by analyzing the magnitude of the interaction forces exerted during the procedure. Both metrics are compared with reference values reported in the literature for manual execution. Finally, system usability is evaluated through subjective measures, including the NASA-Task Load Index (NASA-TLX) questionnaire for perceived workload, the System Usability Scale (SUS) questionnaire for usability, and additional 5-point Likert-scale questions to further characterize the operator interaction. The complete questionnaire is reported in the supplementary materials.
V-B Results
All participants were able to successfully complete the entire experimental procedure. Regarding system efficacy, the mean cleanness score per participant, averaged over the three teeth, ranges from a minimum of $3.66$ to a maximum of $4.28$, indicating a generally high and homogeneous level of task execution. No statistically significant differences between groups are observed, as confirmed by the Kruskal–Wallis test conducted at a significance level of $p=0.05$.
Regarding execution time, the duration of the procedure per tooth showcases substantial variability, ranging from $85\text{\,}\mathrm{s}$ to $212\text{\,}\mathrm{s}$. The median execution times are $153.5\text{\,}\mathrm{s}$ for Group 1, $162.0\text{\,}\mathrm{s}$ for Group 2, and $147.5\text{\,}\mathrm{s}$ for Group 3. Reference values reported in [49] for manual performance of the same procedure indicate a median execution time of $115.5\text{\,}\mathrm{s}$. Although the teleoperated procedure is therefore slower than manual execution, participants exhibited an average reduction of $20.4\%$ in completion time in the second trial, followed by a further $23.2\%$ reduction in the third trial. This trend reflects a clear learning-curve effect, indicating that operators progressively familiarize themselves with the system and improve their ability to control the remote tool. This suggests that, with more extensive training, teleoperated performance could gradually approach manual execution levels.
In terms of interaction forces, during the experimental trial, the force magnitude results in an overall mean of $1.78\pm 0.19\,\mathrm{N}$ for Group 1, $1.86\pm 0.34\,\mathrm{N}$ for Group 2, and $1.96\pm 0.32\,\mathrm{N}$ for Group 3, with average maximum peaks of $9.44\pm 3.09\,\mathrm{N}$, $9.80\pm 4.7\,\mathrm{N}$, and $11.47\pm 4.41\,\mathrm{N}$, respectively. When comparing groups, Kruskal-Wallis tests find no statistically significant differences between experience levels. The comparison with reference values reported in the literature for dentists performing the procedure free-hand highlights a clear discrepancy. In particular, [57] reports an average interaction force of $0.28\pm$0.33\text{\,}\mathrm{N}$\$, with average maximum peaks of approximately $0.60\text{\,}\mathrm{N}$. These values are close to the tissue safety thresholds reported in the literature for dental procedures, commonly identified at around $0.5\text{\,}\mathrm{N}$ [17]. Notably, these reference values are also compatible with those measured in a dedicated free-hand SCP procedure we conduct on the dental model mounted on the force/torque sensor to directly measure interaction forces. Further details on this experiment are provided in the supplementary materials. This difference reflects an inherent limitation of the system, which is primarily attributable to the mechanical characteristics of the adopted industrial robot and to the resolution of the force/torque sensor, not specifically designed to accurately manage the low-magnitude forces typical of dental interactions. Despite the higher measured forces, no superficial damage was observed on the dental model, suggesting that the system remains safe under the adopted experimental conditions. However, these results should be regarded as scalable, as the interaction forces are strongly influenced by the employed hardware setup and are expected to decrease when adopting an engineered system tailored to the target application, particularly in terms of mechanical compliance and force/torque sensing resolution.
The results of the Kruskal–Wallis tests on questionnaire responses are reported in Table I. No statistically significant differences emerge across any NASA-TLX dimension, SUS score, or 5-point Likert item ($p>0.05$), indicating that subjective evaluations are consistent across experience levels. Across all groups, the NASA-TLX results indicate a consistently moderate perceived workload, with contained mental, physical, and temporal demands, limited variability in performance ratings, and generally low effort and frustration. Overall, the task is perceived as manageable and not overly demanding. System usability is evaluated positively, with SUS scores falling within the good-usability range [8], confirming that the system is considered intuitive and easy to use. Likert-scale evaluations further highlight high ratings for interaction-related aspects, including motion and force correspondence, first contact and prolonged interaction perception, and the naturalness of the haptic feedback (mean $\in[3.43,4.33]$), suggesting coherent and supportive sensory feedback. Ergonomics is also positively evaluated (mean $\in[3.71,4.17]$), exceeding the manual reference condition of $3$ and indicating improved ergonomic and postural comfort during teleoperation. Fidelity achieves only moderately positive scores, likely due to limitations in depth perception (mean $\in[1.71,2.33]$), preventing accurate situational awareness.
Group 1 Group 2 Group 3 K–W test Mean $\pm$ STD Mean $\pm$ STD Mean $\pm$ STD $\chi^{2}$ p NASA-TLX Mental Demand $30.00\pm 27.57$ $51.43\pm 15.74$ $36.67\pm 23.38$ 2.91 0.23 Physical Demand $26.67\pm 24.22$ $20.00\pm 11.55$ $16.67\pm 19.66$ 0.70 0.70 Temporal Demand $36.67\pm 23.38$ $42.86\pm 13.80$ $36.67\pm 15.06$ 0.52 0.77 Performance $43.33\pm 8.17$ $7.14\pm 13.80 33$$3.33\pm 10.33$ 2.42 0.30 Effort $16.67\pm 19.66$ $8.57\pm 10.69$ $23.33\pm 15.06$ 2.79 0.25 Frustration $30.00\pm 27.57$ $28.57\pm 27.95$ $33.33\pm 16.33$ 0.11 0.94 Overall $26.19\pm 18.65$ $26.94\pm 13.36$ $25.71\pm 14.26$ 2.00 0.37 SUS $77.92\pm 13.46$ $81.07\pm 9.11$ $78.75\pm 10.34$ 0.51 0.78 5-point Likert Scale Ergonomics $3.83\pm 0.75$ $3.71\pm 0.76$ $4.17\pm 0.98$ 0.95 0.62 Fidelity $3.17\pm 1.33$ $00\pm 0.58 3.3$$33\pm 0.52$ 0.87 0.65 Motion correspondence $4.17\pm 0.41$ $4.00\pm 0.82$ $4.33\pm 0.82$ 0.78 0.68 Force correspondence $4.00\pm 0.89$ $4.14\pm 0.38$ $4.17\pm 0.75$ 0.17 0.92 Depth perception $2.17\pm 0.98$ $1.71\pm 0.76$ $2.33\pm 0.52$ 2.19 0.33 Interaction events $3.83\pm 0.98$ $3.43\pm 0.98$ $4.00\pm 1.10$ 1.30 0.52 Interaction effects $3.83\pm 0.98$ $3.43\pm 0.79$ $4.17\pm 0.75$ 2.73 0.26 Feedback naturalness $4.00\pm 0.89$ $3.71\pm 0.95$ $3.83\pm 1.17$ 0.33 0.85
TABLE I: Summary of the NASTA-TLX and 5-point Likert scale questionnaire responses and SUS scores.
VI Conclusions and Future Perspectives